# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

### v3

@@
Abstract.
Recent advancements in general-purpose AI have highlighted the importance of guiding AI systems towards the intended goals, ethical principles, and values of individuals and groups, a concept broadly recognized as alignment. However, the lack of clarified definitions and scopes of human-AI alignment poses a significant obstacle, hampering collaborative efforts across research domains to achieve this alignment. In particular, ML- and philosophy-oriented alignment research often views AI alignment as a static, unidirectional process (i.e., aiming to ensure that AI systems’ objectives match humans) rather than an ongoing, mutual alignment problem (Wikipedia, 2024). This perspective largely neglects the long-term interaction and dynamic changes of alignment. To understand these gaps, we introduce a systematic review of over 400 papers published between 2019 and January 2024, spanning multiple domains such as Human-Computer Interaction (HCI), Natural Language Processing (NLP), Machine Learning (ML). We characterize, define and scope human-AI alignment. From this, we present a conceptual framework of “Bidirectional Human-AI Alignment” to organize the literature from a human-centered perspective. This framework encompasses both 1) conventional studies of aligning AI to humans that ensures AI produces the intended outcomes determined by humans, and 2) a proposed concept of aligning humans to AI, which aims to help individuals and society adjust to AI advancements both cognitively and behaviorally. Additionally, we articulate the key findings derived from literature analysis, including literature gaps and trends, human values, and interaction techniques. To pave the way for future studies, we envision three key challenges and give recommendations for future research.
Human-AI Alignment, Human-AI Interaction
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Manuscript; submitted to ACM; 2024ccs: Human-centered computing HCI theory, concepts and models
1. Introduction
Artificial Intelligence (AI) has advanced significantly, especially with the advent of general-purpose generative AI, demonstrating unprecedented capabilities in solving a wide range of complicated and challenging problems, such as reasoning, generation, language understanding, and more (Ouyang et al., 2022). However, as AI becomes increasingly powerful and integrated into daily life, society is confronted with an array of risks to both individuals and to broader society (Morris et al., 2024). For example, text-to-image generative models were found to amplify stereotypes about race and gender (blo, 2023), and biased algorithms in hiring processes were found to perpetuate discrimination (Park et al., 2021a). These risks illuminate foundational questions around how and which values are embedded in AI models that drive decisions within real-world contexts.
[…]
@@ 1. Introduction
Figure 1. The overview of the Bidirectional Human-AI Alignment framework. Our framework encompasses both Refer to caption conventional studies of “Align AI to Humans” in AI development that ensures AI produces the intended outcomes determined by humans, and  Refer to caption a novel concept of “Align Humans to AI” , which aims to help humans and society to better understand, critique, collaborate, and adapt to transformative AI advancements. Note that “Humans” refers to both AI users and those who do not interact with AI but may be impacted by AI systems. We further identify four key research questions (i.e., RQ1-RQ4) to facilitate this holistic loop of “bidirectional human-AI alignment”, and organize the literature that can potentially address RQ1-RQ4 in Section 4.
This paper conducts a systematic literature review of over 400 human-AI alignment papers across multiple disciplines, based on PRISMA guidelines (Page et al., 2021; Stefanidi et al., 2023). The literature review corpus includes papers drawn from the the Human-Computer Interaction (HCI), Natural Language Processing (NLP) and Machine Learning (ML) domains that were published between the advent of general-purpose generative AI to present, i.e., primarily between January 2019 and January 2024. We take an interdisciplinary approach to human-AI alignment (Gilbert, 2024), drawing from theories and empirical studies across the aforementioned domains. To round out our approach, we composed our research team of members from the domains of HCI, NLP, ML, Data Science, Computational Social Science, and Cognitive Science.
    Definition. Bidirectional Human-AI Alignment is a comprehensive framework that encompasses two interconnected alignment processes: ‘Aligning AI to Humans’ and ‘Aligning Humans to AI’. The former focuses on integrating human specifications to train, steer, and customize AI, while the latter investigates human cognitive and behavioral adaptations to AI, which supports humans in understanding, critiquing, collaborating with, and adapting to AI advancements.
Paper Organization and Key Findings. We provide clarified definitions and scopes related to human-AI alignment, including “what is the goal of alignment?”, “with whom to align?”, and “what values should be aligned with?”(Section 3). We then present a conceptual framework of Bidirectional Human-AI Alignment from a long-term and dynamic perspective, encompassing both “Align AI to Humans” and “Align Humans to AI” (see Figure 1). Furthermore, we identify four key research questions (RQs) in the Bidirectional Human-AI Alignment framework and provide a structured way to organize the existing research literature to address these questions in Section 4. The resulting structured topologies in Figure 2 and 3 aim to provide a shared vocabulary that can help streamline communication and collaboration between alignment researchers in different disciplines. Furthermore, through iterative paper coding and literature analysis, we derive insights including meta analysis on gaps and trends in existing alignment research, findings of human values, and the interaction techniques for human-AI alignment (Section 5). To pave the way for future studies, we further envision three future directions in Figure 6 and Section 6. Our study highlights several key findings. For instance, many essential human values outlined in Table 2, such as loyalty and environmental protection, are crucial for alignment but are often overlooked in current research. Additionally, many topics within the bidirectional alignment framework, including leveraging implicit human feedback to align AI models and auditing AI for various social values, are important but underexplored.
Operationalizing the Framework. To help AI practitioners (e.g., developers, researchers, user experience designers) who are interested in this field to leverage the bidirectional approaches, we highlight some systematic resources about human-AI alignment provided both within this survey and on companion Github repository:
  • A Github repository of “Bidirectional Human-AI Alignment” maintaining latest reading lists and relevant updates.
  • Fig. 1: An overview of the bidirectional human-AI alignment framework with formal definition.
  • Table 1: Typical alignment goals, their definitions, and limitations.
  • Table 2: A comprehensive list of basic human values, their relationships, and related alignment papers.
  • Fig. 2 and 3: Fine-grained topologies and papers of “Aligning AI to Humans” & “Aligning Humans to AI” research.
  • Fig. 4: An illustration of trends and gaps of alignment research dimension distribution.
  • Fig. 5: A summary of common interactive techniques to specify human values and exemplary papers.
2. Challenges in Achieving Alignment
[…]
Challenge 1: Outer and Inner Alignment. In the context of “intelligent agents,” until now, AI alignment research has aimed to ensure that any AI systems that would be set free to make decisions on our behalf would act appropriately and reduce unintended consequences (Wooldridge, 1999; Stuurman and Wijnands, 2001; Russell and Norvig, 2016). At the near-term stage, aligning AI involves two main challenges: carefully specifying the purpose of the system (outer alignment i.e., providing well-specified rewards (Ngo et al., 2024)) and ensuring that the system adopts the specification robustly (inner alignment, i.e., ensuring that every action given an agent in a particular state learns desirable internally-represented goals (Ngo et al., 2024)). Significant efforts have been made, for inner alignment, to align AI systems to follow alignment goals of an individual or a group (e.g., instructions, preferences, values, and/or ethical principles) (Ouyang et al., 2022) and to evaluate the performance of alignment (Santurkar et al., 2023). However, for outer alignment, AI designers are still facing difficulties in specifying the full range of desired and undesired alignment goals of humans.
Challenge 2: Specification Gaming. To learn human alignment goals, AI designers typically provide an objective function, instructions, reward function, or feedback to the system, which is often unable to completely specify all important values and constraints that a human intended (Hemphill, 2020). Hence, AI designers resort to easy-to-specify proxy goals such as maximizing the approval of human overseers (Wikipedia, 2024), which results in “specification gaming” (Krakovna et al., 2020) or “reward hacking” (Pan et al., 2022) issues (i.e., AI systems can find loopholes that help them accomplish the specific objective efficiently but in unintended, possibly harmful ways). Additionally, the black-box nature of neural networks brings additional ethical and safety concerns for alignment because humans don’t know about the inner state and the actions AI leveraged to achieve the output. Consequently, AI systems might make “correct” decisions with “incorrect” reasons, which are difficult to discern. Society is already facing these issues, such as data privacy (Tucker et al., 2018), algorithmic bias (Hajian et al., 2016), self-driving car accidents (Bonnefon et al., 2016), and more. As a result, these considerations necessitate considering human-AI interaction in AI alignment for specification and evaluation, ranging from addressing problems around who uses an AI system, with what goals to specify, and if the AI system perform its intended function from the user’s perspective (Weidinger et al., 2023).
Challenge 3: Scalable Oversight. From a long-term perspective, when advanced AI systems become more complex and capable (e.g., AGI (Morris et al., 2024)), it becomes increasingly difficult to align them to human values through human feedback. Evaluating complex AI behaviors applied to increasingly challenging tasks can be slow or infeasible for humans to ensure all sub-steps are aligned with their values (Terry et al., 2023). Therefore, researchers have begun to investigate how to reduce the time and effort for human supervision, and how to assist human supervisors, referred to as Scalable Oversight (Amodei et al., 2016).
Challenge 4: Dynamic Nature. As AI systems become increasingly powerful, the alignment solutions must also adapt dynamically since human values and preferences change as well. As Dautenhahn et al. (2000) posit, AI systems may be neither humane nor desirable if we do not ask questions about the long-term cognitive and social effects of social agent systems (e.g., how will agent technology affect human cognition). All these considerations call for a long-term and dynamic perspective to address human-AI alignment as an ongoing, mutual process with the collective efforts of cross-domain expertise.
Challenge 5: Existential Risk. Further, some AI researchers claim that (Bengio et al., 2024) advanced AI systems will begin to seek power over their environment (e.g., humans) once deployed in real-world settings, as such behavior may not be noticed during training. For example, some language models seek power in text-based social environments by gaining money, resources, or social influence (Pan et al., 2023). Consequently, some hypothesize that future AI, if not properly aligned with human values, could pose an existential risk to humans (Dafoe and Russell, 2016).
Table 1. The Goals of Alignment. We present the six prevailing alignment goals, associating with their Definitions (middle column), Limitations and Risks (right column). We consider Human Values as the main goal of alignment in this work referring to an extensive analysis and arguments in existing studies (Gabriel, 2020; Russell, 2014)
3. Fundamental Definitions and Clarifications
Addressing alignment challenges is a complex and multifaceted process. To develop a comprehensive perspective of the ongoing and mutual process of human-AI alignment, this section begins by introducing core definitions and key components of alignment. Subsequently, we outline our methodology for conducting a systematic literature review and the rationale behind our proposed framework of bidirectional human-AI alignment.
3.1. What is the goal of alignment?
[…]
Table 2. The relations and a fine-grained taxonomy of 69 exemplary human values. We consider 5 high-order value types encompassing 12 motivational value types, indicated by their sources (e.g., individuals, society and interaction). The exemplary values with red dot ([Uncaptioned image]) indicates there are no work in our surveyed papers examining the specific values.
3.2. With whom to align?
Pertinent stakeholders within the AI landscape can be the potential objects for AI to align with, including lay people (e.g., end users of AI systems) (Qi et al., 2023), AI practitioners (e.g., developers, researchers) (Boggust et al., 2022; Anwar et al., 2024; Ngo et al., 2024), organizational entities (e.g., technology firms, professional communities), national/international bodies (e.g., governments, legislative bodies) (Deshpande and Sharp, 2022) and others. Many alignment research papers have focused on general humans without specifying particular groups (Yao et al., 2023b; Liu et al., 2022; Guo et al., 2024). Nevertheless, different groups hold different, sometimes even contrasting, values (Gabriel, 2020). As a consequence, rather than identifying a true moral theory as a one-size-fits-all value, prior alignment research argues to select the appropriate principles for compatible human groups (Sorensen et al., 2024). To this end, pluralistic value alignment, grounded on social choice theory (Arrow, 2012), proposes combining individual views fairly in developing alignment principles as a potential solution (Sorensen et al., 2024). In this work, we also consider values from this pluralistic perspective, where AI should be aligned with pluralistic human individuals and societal groups who would ultimately be impacted by AI.
3.3. What are the values to be aligned with?
[…]
A Comprehensive Taxonomy of Human Values. Nevertheless, this conventional theory was developed without the context of human-AI interaction, which might overlook values that need to be considered for human-AI alignment. Therefore, we used a bottom-up approach to extract all values studied in our collected alignment literature (elaborated in Section A.1), mapped them onto the Schwartz Theory of Basic Values, and supplemented the theory with AI-related structure and content. As a result, we identified the structural relationships among human values (see Figure LABEL:fig:value_structure) and mapped existing literature to a fine-grained taxonomy (see Table 2). As shown in Figure LABEL:fig:value_structure, we supplemented the traditional theory’s four high-order value types (i.e., “Self-Enhancement”, “Openness to Change”, “Conservation”, “Self-Transcendence”) with a novel high-order value type, named “Desired Values for AI Tools” that encompasses two motivational value types (i.e., “Usability” and “Human-Likeness”). We further organize the relationship among these value types along two dimensions (Schwartz, 2012): different resources (i.e., individuals, society and interaction) and different self-intentions (i.e., self-protection against threat and self-expansion and growth). Furthermore, we elaborate the definitions of the 12 motivational value types and their exemplary values by mapping them to relevant human-AI alignment papers from our corpus in Table 2. During the process of mapping, we found: 1) value terms in empirical papers were often named differently (e.g., capability and competence), or check their opposites (e.g., fairness and bias); 2) there are many values not studied in our corpus, i.e., indicated as ([Uncaptioned image]) in the Figure.
    Takeaways & Implications for Alignment Clarifications 1. We propose the “pluralistic human values” as the alignment goal, wherein AI should be aligned with the diverse perspectives of individuals and societal groups who may be directly or indirectly impacted by AI. 2. Many exemplary human values outlined in Table 2, such as loyalty and environmental protection, are important for alignment but are often overlooked or underexplored in current research. 3. The human value system should account for complex application contexts and relationships, such as priorities and conflicts, rather than focusing on an exclusive subset of values.
3.4. Scopes and Key Components in Alignment
This section clarifies the key components we considered in human-AI alignment. Additionally, we define the scopes of each key component as commonly adopted in current research.
  • Humans. Our primary focus is on human individuals, groups, or organizations that will ultimately develop, use, and potentially impact or be influenced by AI systems, as these are the entities with which AI systems should align. We emphasize the importance of considering the pluralistic values of diverse users rather than treating users as a monolithic group.
  • Artificial Intelligence (AI). We focus on AI systems including both domain-specific AI systems that address specific tasks (e.g., reasoning, dialogue) and general-purpose AI models that aim to complete any tasks with performance comparable to a human’s. These AI systems include generative, classification, and regression models, among others. Particularly, we primarily focus on language models as the representative AI models for alignment research, and discuss how the insights from this study can be generalized to other modalities at the section end.
  • Alignment. Our review encompassed all the alignment goals outlined in Section 3.1. Since many studies emphasized the importance of value alignment (Gabriel, 2020; Russell, 2014), we particularly summarized a clarified taxonomy of alignment values and identified the value in each paper (if applicable). Additionally, we focus on analyzing the AI models’ output and generation, but not the neural network’s intermediate representations (McGrath et al., 2023; Park et al., 2023; Xie et al., 2024), for alignment research.
3.5. Systematic Literature Review and Framework Rationale
Achieving the alignment between AI and human values necessitates collaboration among interdisciplinary AI practitioners (e.g., developers and researchers). To comprehensively analyze the research literature pertaining to this ongoing, mutual process of human-AI alignment, we conducted a systematic literature review adhering to the the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023). Our iterative selection process yielded 411 papers from an initial pool of 34,213 publications, spanning the period from January 2019 to January 2024. These papers were sourced from leading venues across multiple AI-related domains, including Natural Language Processing, Human-Computer Interaction, Machine Learning, and more, using well-defined keywords and criteria. A group of 13 interdisciplinary researchers coded the papers. Through iterative discussions, we developed the paper codes and proposed framework (Section 3). Detailed steps and processes are elucidated in Appendix A.
3.6. Rationale Behind the Framework
Our systematic review revealed that current research largely overlooks the nature of long-term dynamic changes and the importance of interaction in human-AI alignment. First, the manifestation of certain AI risks, harms, and the practical quality of AI services primarily emerge through interaction between humans and AI systems (Weidinger et al., 2023). For instance, Shen and Huang (2020) discovered that AI explanations, deemed effective by AI developers and researchers, were found not useful for humans in real-world scenarios of interactive AI error detection. Additionally, the use of AI influences human thinking and behavior, which in turn alters their values when interacting with AI. For instance, Li et al. (2024b) found that incorporating AI assistance in writing workflows changed humans’ writing perceptions, behavior, and performance, creating new requirements for developing the next generation of AI writing assistants (Lee et al., 2024).
In light of these findings, we posit that AI practitioners must expand their focus beyond merely “Aligning AI to Humans” – a paradigm centered on integrating human specifications to train, steer, and customize AI. Equal emphasis must be placed on “Aligning Humans to AI”, a complementary paradigm investigating human cognitive and behavioral adaptation to AI. This latter aspect is crucial in supporting individuals and society at large in understanding, critiquing, collaborating with, and adapting to transformative AI advancements. To encapsulate these comprehensive perspectives, we propose the “Bidirectional Human-AI Alignment” framework, and provide its definition below Figure 1. This holistic framework acknowledges the reciprocal nature of alignment, addressing both the advancements of AI systems to incorporate human values, and the evolving human adaptation of AI systems in real-world practice.
4. Bidirectional Human-AI Alignment Framework
Figure 2. The fine-grained topology of “Align AI to Humans” direction, which studies mechanisms to ensure that AI system’s objectives match those of humans’. The goal is to integrate human specification to train, steer and customize AI systems.
This section introduces the Bidirectional Human-AI Alignment framework, which encompassing two interconnected alignment directions, as shown in Figure 1. The [Uncaptioned image] “Align AI to Humans” direction studies mechanisms to ensure that AI systems’ values match those of humans’ (Section 4.1 and 4.2). We identify two crucial research questions to be addressed in this direction: RQ1. What relevant human values are studied for AI alignment, and how do humans specify these values? and RQ2. How can human values be integrated into the development of AI? In comparison, the [Uncaptioned image] “Align Humans to AI” direction investigates the humans’ cognitive and behavioral adaptation to the AI advancement (Section 4.3 and 4.4). We recognize two imperative research questions along this direction: RQ3. How might humans learn to perceive, explain, and critique AI? and RQ4. How do humans and society make behavioral changes and react to AI advancement?. We address the four questions by elucidating the fine-grained typologies and literature solutions derived from our systematic review, as delineated in Figure 2 and Figure 3, respectively.
DIRECTION-I:  [Uncaptioned image] ALIGN AI to HUMANS
This direction delineates alignment research from AI-centered perspective (e.g., ML/NLP domains) and provides AI developers and researchers with approaches for two main challenges: carefully specifying the values of the system, and ensuring that system adopts the specification robustly (Wikipedia, 2024; Ngo et al., 2024). Therefore, as shown in Figure 2, we explore the two core research questions in this direction as: RQ1. Human Values and Specifications (Section 4.1) and RQ2. Integrating Human Specifications into AI (Section 4.2)
4.1. Align AI to Humans: [Uncaptioned image] Human Values and Specifications
RQ1: What relevant human values are studied for AI alignment, and how do humans specify these values? We structure existing literature to address this research question by answering the following “Sub-Research Questions”: What values have been aligned by AI? (Section 4.1.1) and then exploring How humans could interactively specify values in AI development? (Section 4.1.2). As shown at the top of Figure 2, particularly, we articulate the “Dimensions“ we summarized to answer each of these sub-research questions, and provide “Codes“ associated with “Example Papers” that have been studied in each dimension.

### v4

# Position: Towards Bidirectional Human-AI Alignment

@@
Abstract
Recent advances in general-purpose AI underscore the urgent need to align AI systems with human goals and values. Yet, the lack of a clear, shared understanding of what constitutes "alignment" limits meaningful progress and cross-disciplinary collaboration. In this position paper, we argue that the research community should explicitly define and critically reflect on "alignment" to account for the bidirectional and dynamic relationship between humans and AI. Through a systematic review of over 400 papers spanning HCI, NLP, ML, and more, we examine how alignment is currently defined and operationalized. Building on this analysis, we introduce the Bidirectional Human-AI Alignment framework, which not only incorporates traditional efforts to align AI with human values but also introduces the critical, underexplored dimension of aligning humans with AI – supporting cognitive, behavioral, and societal adaptation to rapidly advancing AI technologies. Our findings reveal significant gaps in current literature, especially in long-term interaction design, human value modeling, and mutual understanding. We conclude with three central challenges and actionable recommendations to guide future research toward more nuanced, reciprocal, and human-AI alignment approaches.
1 Introduction
Artificial Intelligence (AI), particularly generative AI, has demonstrated remarkable capabilities in reasoning, language understanding, problem solving, and more ouyang2022training . However, its increasing integration into society raises significant risks, such as amplifying biases in hiring bloomberg or perpetuating stereotypes in text-to-image models park2021human . These concerns highlight the urgent need to align these systems with values, ethical principles, and the goals of individuals and society at large. This need, commonly referred to as “AI alignment,” wiki_AI_alignment ; terry2023ai is crucial for ensuring that AI systems function in a manner that is not only effective but also consistent with human values, minimizing harm and maximizing societal benefits. Yet, key challenges remain:
Challenge 1: Specification Gaming. AI designers often define objectives or feedback to align systems with human goals, but these rarely capture all intended values hemphill2020human . This leads to reliance on proxies like human approval wiki_AI_alignment , enabling specification gaming krakovna2020specification ; pan2022effects , where AI makes seemingly “correct” decisions for the wrong, opaque reasons tucker2018privacy ; hajian2016algorithmic ; bonnefon2016social .
Challenge 2: Scalable Oversight. As AI systems become more complex—potentially reaching AGI morris2024levels —aligning them through human feedback grows harder. Evaluating their behavior is often slow or infeasible terry2023ai , prompting research into reducing supervision burdens and enhancing human oversight, a challenge known as Scalable Oversight amodei2016concrete .
Challenge 3: Dynamic Nature. As AI advances, alignment must adapt to evolving human values. Without considering long-term cognitive and social impacts, AI may become neither humane nor desirable dautenhahn2000living . This needs a dynamic, ongoing alignment process with cross-disciplinary collaboration.
@@ 1 Introduction
Figure 1: The overview of the Bidirectional Human-AI Alignment framework. Our framework encompasses both Refer to caption conventional studies which focus on “Align AI with Humans” and  Refer to caption “Align Humans with AI”. We further identify four key Research Questions to facilitate this holistic loop of “bidirectional human-AI alignment”, and provide answers that can potentially address RQ1-RQ4 in Section 3.
Traditionally, AI alignment has been approached as a static, one-way process focused on shaping AI to achieve desired outcomes and avoid harm goyal2024designing ; ouyang2022training ; santurkar2023whose . However, this unidirectional view is increasingly insufficient as AI systems become more integrated into daily life and assume complex decision-making roles carroll2024ai . Their interactions with humans create evolving feedback loops that influence both AI behavior and human responses shen2020useful ; li2024value ; lee2024design , highlighting the need for a more dynamic and reciprocal understanding of alignment carroll2024ai .
In this position paper, we argue that it is critical for the research community to explicitly reflect on what we mean by “alignment” and to take into account the bidirectional, dynamic interactions between humans and AI to achieve responsible and safe AI systems.
Through a systematic review of over 400 papers across HCI, NLP, ML, and related fields, we examine how alignment is currently defined and implemented. Based on this analysis, we propose the Bidirectional Human-AI Alignment (in Figure 1) framework. It extends the traditional focus on “Aligning AI to Humans”—integrating human input to train, steer, and customize AI—by introducing the equally vital yet underexplored direction of “Aligning Humans to AI”, which emphasizes cognitive, behavioral, and societal adaptation to rapid AI advancement.
Our findings reveal key gaps in existing research, particularly in human value modeling, oversight of model inference, critical evaluation of AI’s embedded values, and its broader societal impact. We conclude by outlining near- and long-term risks and opportunities, offering actionable recommendations to advance more reciprocal, adaptive, and nuanced approaches to human-AI alignment.
2 Defining Alignment: Fundamentals
Building on our analysis of systematic review (see details in Appendix 8), we explicitly identify the key definitions in alignment research and formally propose “Bidirectional Human-AI Alignment”.
Goals. AI alignment research proposes multiple alignment goals park2023generative ; xie2024can , such as intentions ouyang2022training ; anwar2024foundational , preferences bansal2023peering ; verma2023hindsight , instructions bai2022constitutional ; liu2023makes , and values sorensen2024roadmap ; gabriel2020artificial . But these terms are often used interchangeably without clear distinctions. Philosophical analysis suggests human values (moral beliefs/principles) are the most suitable alignment goal, as they ensure AI acts ethically while minimizing risks gabriel2020artificial ; stuart2014whitepaper . Though trade-offs exist, this work adopts "human values" as the alignment objective, meaning AI should behave as individuals or society morally expect (See details in Table 2).
Align with Whom. AI alignment involves multiple stakeholders, such as end users qi2023fine , AI practitioners boggust2022shared ; anwar2024foundational ; ngo2022alignment , and organizations deshpande2022responsible . Many studies reference “general humans” without accounting for group differences, despite the fact that stakeholders often hold conflicting values gabriel2020artificial . Rather than pursuing a universal moral theory, prior work advocates aligning AI with principles tailored to compatible human groups sorensen2024roadmap . Pluralistic value alignment, grounded in social choice theory arrow2012social ; sorensen2024roadmap , provides a framework for integrating diverse perspectives sorensen2024roadmap . We adopt this view, recognizing that aligning with affected individuals and groups entails ongoing challenges.
Align with What. While prior studies have sought to align AI with human values, these values are often vaguely or inconsistently defined. To address this, we reviewed several prominent value theories, including Moral Foundations Theory graham2013moral and Social Norms and Ethics forbes2020social , drawing from psychology and social science. We selected the Schwartz Theory of Basic Values schwartz1994there ; schwartz2012overview for its demonstrated cross-cultural applicability, relevance across contexts (individuals, interactions, groups), and frequent adoption in NLP research kang2023values ; kiesel2022identifying . This framework consists of eleven types of universal values, and defines values as beliefs about desirable end states or behaviors that transcend specific situations and guide evaluation and decision-making.
Definition: Bidirectional Human-AI Alignment is a comprehensive framework that encompasses two interconnected alignment processes: ‘Aligning AI with Humans’ and ‘Aligning Humans with AI’. The former focuses on integrating human specifications into training, steering, and customizing AI. The latter supports human agency, empowering people to think critically when using AI, collaborate effectively with it, and adapt societal approaches to maximize its benefits for humanity.
3 Bidirectional Human-AI Alignment Framework
Table 1: The fine-grained typology of two directions in the Bidirectional Human-AI Alignment.
Research Question Sub-Research Question Dimensions References